OpenStudy (anonymous):

prove : A + B = π/4 if tanA = m/m+1 , tanB = 1/2m+1 (m ≥ 0) pls help....

OpenStudy (anonymous):

tan(A+B) = tanA + tanB / 1 - tanAtanB = m/m+1 + 1/2m+1 / 1 - (m/m+1* 1/2m+1) =( m(2m+1) + m + 1)/ (m+1)(2m+1) * (2m+1)(m+1)/(2m+1)(m+1) - m = 2m^2 + 2m +1 /2m^2 + 2m + 1 = 1 so tan(A+B) = 1 and A+B = 45degrees = pi/4 radians

OpenStudy (anonymous):

thanks a lot.. :)